Utica Company's net accounts receivable was $250,000 at December 31, 2016, and $300,000 at December 31, 2017. Net cash sales for 2017 were $100,000. The accounts receivable turnover for 2017 was 5.0, which was computed from net credit sales for the year.

What was Utica's total net sales for 2017?

(This is a previously posted question that was answered incorrectly).
Answer #1

Receivables turnover = Net credit sales / average inventory

average inventory = (250,000 + 300,000) / 2 = 275,000

average inventory = 275,000

5 = Net credit sales / 275,000

Net credit sales = $1,375,000

Total net sales = Cash sales + credit sales

Total net sales = 100,000 + 1,375,000

Total net sales = $1,475,000